OpenStudy (anonymous):

i am having trouble with a back tracing equation. 4-2a=2

OpenStudy (anonymous):

tracking not tracing sorry

OpenStudy (anonymous):

umm, i don't know about back tracking but a = 1

OpenStudy (anonymous):

thanks we had the answer but needed to show how to solve it.

OpenStudy (anonymous):

4-2a=2 -2a = -2 subtract both sides with 2 a = 1 divide both sides with -2

OpenStudy (anonymous):

I need to show the working out from the answer back to the beginning

jhonyy9 (jhonyy9):

4-2=2a 2=2a 2/2 =a a=1

OpenStudy (anonymous):

[(4-2a)/2] - 1 = 0 well, you can just multiply or divide numbers randomly to make it more complicated

OpenStudy (anonymous):

I think op just simply means 4 -2*1 = 2......:-) Maybe..